OpenStudy (anonymous):

Shirley purchased a plot of land for $19,500. The land appreciates about 3.9% each year. What is the value of the land after five years? $23,610.89 $19,818.94 $15,982.75 $23,302.50

OpenStudy (anonymous):

First you find how much is appreciating in a year: \[\frac{ 3.9 }{ 100 }*19.500=0.7605\] Then you multiply this value by 5: \[0.7605*5=3.8025\] And add it to the initial VALUE: \[19.500+3.8025=23.302.50\]

OpenStudy (anonymous):

23,302.50

OpenStudy (anonymous):

Thank you SOO much. In step 1, how do you know where to place the decimal in amounts like 19,500? Is it always three numbers to the left?

OpenStudy (anonymous):

Yeah it is always three numbers to the left but... If your last group is formed by 0's as in case "23.302.500" you can obmit the last zero or the last 2.

OpenStudy (anonymous):

Oh, Alright. Thanks (: